Line of symmetry for y=5x^2+4x-2
Answer by stanbon(75887)

Line of symmetry for y=5x^2+4x-2
--------
Line of symmetry is x = -b/(2a)
----
Ans:: x = -4/(2*5) = -4/10 = -2/5
------------------
Cheers,
Stan H.
--------------